I asked this a few days ago but didn't get an answer. Can someone help me with this please?




When using the quadratic formula to solve a quadratic equation ax2 + bx + c = 0, the discriminant is b2 - 4ac. This discriminant can be positive, zero, or negative. (When the discriminate is negative, then we have the square root of a negative number. This is called an imaginary number, sqrt(-1) = i. )
Explain what the value of the discriminant means to the graph of y = ax2 + bx + c. Hint: Chose values of a, b and c to create a particular discriminant. Then, graph the corresponding equation. In the real world, where might these imaginary numbers be used?

Answer by ilana(307)

When the descriminant is positive, the parabola is either like a U with the vertex below the x-axis, or like an upside-down U with the vertex above the x-axis (so either way, the parabola intersects the x-axis at 2 points). When the descriminant is 0, the vertex is on the x-axis (like y=x^2), and when the descriminant is negative it never touches the x-axis. So it either starts above and continues up or starts below and continues down (like y=x^2+3 or y=-x^2-3).
I believe that's part of what the question is looking for.